# OmniBal: Towards Fast Instruct-tuning for Vision-Language Models via Omniverse Computation Balance

### v3

@@ 4 Method
This section presents our computation-balanced framework OmniBal for training large vision-language models. To address an imbalanced computational load across devices, we first manage the large data input variations, which is the most fundamental issue in the computation imbalance problem. This enables balanced model partitioning. Finally, the re-computation strategy for each partition is optimized. Figure 3 shows our training pipeline.
@@ 4.1 Balanced Dynamic Mini-Batch
For instruct-tuning VLMs, each training sample contains various images and texts, resulting in non-fixed input sizes. We evaluate data imbalance from two perspectives: within-device samples and cross-device mini-batches.
@@
Dist ratio (cross-device): Even after padding, the sizes of mini-batches on different devices may vary, leading to different input scales across devices. The distribution ratio is calculated as follows:
@@
The sampling and filtering stages alternately are repeated for a maximum of $T$ times. The candidate set is acquired $\mathcal{P}$ each time, which includes more valid sample groups $\mathcal{G}$. Meanwhile, we have the updated dataset $\mathcal{D}$ consisting of unselected samples, which is used for sampling and filtering in the next iteration. To ensure that the mini-batches constructed by the ISF method achieve lower Pad Ratio and Dist ratio, appropriate values for $Q_{v}$ and text $Q_{t}$ need to be determined. The optimal values for $Q_{v}$ and $Q_{t}$ vary across different datasets. In practice, A statistical approach described in Section LABEL:sec:impl_detail is used to determine these values.
@@ 4.2 Balanced Model Partitioning
4.3 Balanced Adaptive Re-Computation
We find that heterogeneous architectures have different memory requirements. For example, the vision model in InternVL-Chat-1.5 requires more GPU memory than the language model under the same computational load. Therefore, it is necessary to analyze the memory requirements of each layer in the vision and language models individually and adaptively determine the optimal re-computation strategy for each layer. Specifically, we start by profiling to determine the memory requirements of each layer. Based on the available memory of each device, we then determine the optimal re-computation configurations in pipeline stage $S_{i}$.
Here, we provide a detailed introduction to the Balanced Adaptive Re-Computation Strategy. In this context, $Q_{v}$ and $Q_{t}$ represent the inputs for Vision Transformer (VIT) and Large Language Model (LLM) respectively. $M_{r}$ denotes the remaining GPU memory at the current stage, while $M_{t}$ and $M_{v}$ indicate the GPU memory saved by each transformer layer of the LLM and VIT when enabling re-computation.
5 Experiments
In this section, The models and datasets are introduced. Then, we demonstrate the acceleration compared to current state-of-the-art VLMs. Subsequently, a detailed comparison of each component proposed in our method is presented, highlighting its specific contribution to training acceleration. Finally, extensive experimental analysis is conducted.
@@ 5.1 Experimental setup
Implementation Details: We determine $Q_{v}$ and $Q_{t}$ using dataset statistics. The total text token lengths and image count are used to compute the average tokens per image. $Q_{t}$ is set to the longest text token length, and the text-to-image ratio determines $Q_{v}$. For images, $Q^{\prime}_{v}=Q_{v}$, and for text, $Q^{\prime}_{t}=Q_{t}-128$. In the InternVL-Chat-1.2 dataset, $Q_{t}=4K$, $Q_{v}=9$, with each image processed into 1K tokens for VIT.
Choice of $Q_{v},Q_{t}$. Here we provide some ablation experiments for choosing $Q_{v}$ and $Q_{t}$: $Q_{v}$ represents the threshold for the number of input images, while $Q_{t}$ represents the threshold for text length. It is challenging to maintain an exact number of images and text length, so we relax these conditions to allow for approximation.
Partition computation The computational overhead is minimal. In fact, establishing the basic anchor strategy involves profiling the forward time and activation values of different layers. In our experiment, only five steps are needed, resulting in very little computational overhead. The anchor may not be the direct optimal solution. We empirically find that the final optimal solution is generally within $r\leq 3$. Cost of Calculating the Optimal Solution: After obtaining the anchor, we select the top K partitions (10-15) near the anchor for speed measurement, as described in Section 4.2. Each speed measurement requires only 3 steps, making the cost relatively small compared to the entire training process.
@@
Table 6: Results of $Q_{v},Q_{t}$
@@ 5.2 Main Results
5.3 Ablation Analysis
In this section, ablation experiments on each component of our method are conducted, using InternVL-Chat-1.5 as the baseline model with a 3D parallel Megatron-DeepSpeed backend. Table 7 illustrates the impact of each component. The baseline model experiences a considerable slowdown in training speed due to computational imbalance, necessitating a total of 61.8 GPU days. By achieving data balance, GPU days are reduced to 51.9. Data balance allows us to achieve a more balanced model partition, reducing the training time. Finally, optimizing memory with an adaptive re-computation strategy reduces GPU days to 21.3. These results demonstrate that a holistic balance encompassing data, model, and memory is crucial for efficient VLM training. Below we provide a detailed analysis of each component.
The Importance of Data Balance: In Table 3, we investigate the importance of data balance in large-scale distributed training by comparing four methods: (1) Baseline: Randomly combining data into a mini-batch with padding aligned to the longest input within mini-batches (2) Length-Group: Combining samples with similar text and image sizes into a mini-batch to minimize padding within mini-batch. (3) Device-Group: Grouping samples with similar input sizes across devices to minimize idle times. (4) Balanced Dynamic Mini-batch: Using ISF to construct balanced mini-batches within mini-batches and cross devices.
The Importance of Model Balance: Table 4 examines balanced model partitioning, focusing on partition strategies for pipeline parallelism. For LLM training, common methods include (1) parameter-based and (2) layer-based, (3) profile-based methods such as DreamPipe (Narayanan et al., 2019) estimate the computation time for each layer and use this information to partition the model effectively. Additionally, (4) our search-based Balanced Model Partition method finds the optimal partition strategy from a set of candidates. As shown in Table 4, (1) Parameter-based and (2) layer-based methods split the model’s parameters or layers across devices, achieving low variation in VAR(param) and VAR(num_layer). However, they still show high variation in forward time VAR(fwd_time), leading to computational inefficiencies in the pipeline. (3) The profile-based method ensures the optimal VAR(fwd_time). However, this partitioning occurs before the vision model’s token sub-sampling operation, increasing communication overhead and affecting training speed. (4) Our proposed BMP method explores a high-quality partition strategy space to identify the optimal strategy, achieving the best results in 29.0 GPU days.
The Importance of Memory Balance: In Table 5, we examine the significance of memory balance. In the baseline model, varying input sizes for vision (4K–20K tokens) and language (1K–16K tokens) lead to varying GPU memory usage. Despite aggressive re-computation, the remaining memory can drop to 7.3 G. ISF and BMP improve training speed by controlling computational load across devices. However, memory demands still varied, e.g., GPUs 1, and 2 having more remaining memory. ISF further speeds up training by adjusting the re-computation strategy to fully utilize the remaining memory, reducing GPU days to 21.3.
5.4 Convergence of ISF
5.5 Generalization Capability
Different Models: we verify our methods on different vision models like EVA-CLIP (Sun et al., 2023) and different language models like Llama3 (AI, 2024), Yi-34B (NousResearch, 2023), and the Qwen1.5-110B (Bai et al., 2023a). As shown in Table 8, our approach significantly reduces the required GPU days for model training, achieving nearly a 2$\times$ speedup across models of various sizes.
Results on Different Dynamic High-Resolution Setting: to validate the effectiveness of our method, we test it under various high-resolution settings. Our approach consistently demonstrates low Dist Ratio and strong acceleration across all configurations as shown in Table 9, significantly improving training speed under different settings.
@@ 5.5 Generalization Capability
Results on different datasets: we consistently achieved a low Dist Ratio across the LLava-665K, InternVL-1.2M, and LCS558K datasets, as demonstrated in Table 11. Additionally, our approach significantly enhanced training speed.
Table 11: Results on different datasets
Dataset Dist Ratio GPU Days
VIT LLM
LLava-665K 0.02 0.145 43.3$\rightarrow$ 12.4 (3.5$\times$)
InternVL-1.2M 0.02 0.14 61.8$\rightarrow$ 21.3 (2.9$\times$)
LCS-558K 0.001 0.029 23.8$\rightarrow$ 7.5 ( 3.2$\times$)
Results on Different Resolutions: we further test our method with different image resolution inputs. As shown in Table 12, our method consistently delivers low Dist Ratio and highly satisfactory acceleration results across varying image resolutions, demonstrating its effectiveness in improving training efficiency.
@@
Results on Open-source LLava-1.6: we also validate our method using another popular open-source model, LLava-1.6, with the DeepSpeed backend, as shown in Table 13. For the DeepSpeed backend, we employ only our balanced dynamic mini-batch strategy. In the case of the open-source LLava model, while its ViT component is relatively small and the imbalance occurs primarily at the data level, we still achieved a notable overall speedup. Although the speedup ratio is smaller compared to other models, our method delivered a 30% improvement in performance.
@@
Results on Pretrain Setting: we evaluate our method in other tasks, such as pre-training tasks. In Pre-training, we train both the Vision Transformer (ViT) and MLP components for models ranging from 6B to 20B. However, for larger models, such as 6B-34B and 6B-70B, we focus solely on training the MLP component. Across all configurations, we observe consistent performance improvements shown in Table 10, particularly with the largest model, where GPU days are significantly reduced from 16.8 to 9.6, demonstrating enhanced training efficiency.
Results on Qwen2-VL Pre-Processing Strategy: Qwen2-VL is a recent, highly regarded open-source project that provides strong support for dynamic image input. Consequently, we adopt the pre-processing strategy of Qwen2-VL to validate our method. As shown in Table  14, our approach demonstrates a substantial acceleration effect (approximately 1.9$\times$) when applied to the Qwen2-VL strategy, significantly reducing both the padding ratio and dist ratio.
@@
Different Hardware Results: we test our method on various hardware platforms with different GPUs (e.g., A100, H100) and network bandwidths. The experiments in Table  15 confirmed consistent performance improvements across all platforms.
Table 15: Results on Different Hardware. IB indicates network bandwidths
@@
Large-Scale Results: to validate the effectiveness of our method, we conduct a study using larger-scale models and a greater number of GPUs. As shown in the Tabel 16, our method achieves a speedup ratio exceeding 2.0 across varying GPU configurations. Moreover, the results demonstrate that our approach maintains a more favorable linear speedup (85% $\rightarrow$ 95%) as GPUs increase.

### v4

@@ 4 Method
This section presents our computation-balanced framework, OmniBal, for training large vision-language models. To address an imbalanced computational load across devices, we first manage the large data input variations, which is the most fundamental issue in the computation imbalance problem. This enables balanced model partitioning. Finally, the re-computation strategy for each partition is optimized. Appendix A shows our training pipeline.
@@ 4.1 Balanced Dynamic Mini-Batch
For instruction-tuning VLMs, each training sample contains various images and texts, resulting in non-fixed input sizes. We evaluate data imbalance from two perspectives: within-device samples and cross-device mini-batches.
@@
Distribution Ratio (cross-device): Even after padding, the sizes of mini-batches on different devices may vary, leading to different input scales across devices. The distribution ratio is calculated as follows:
@@
The sampling and filtering stages are alternately repeated for a maximum of $T$ times. The candidate set is acquired $\mathcal{P}$ each time, which includes more valid sample groups $\mathcal{G}$. Meanwhile, we have the updated dataset $\mathcal{D}$ consisting of unselected samples, which is used for sampling and filtering in the next iteration. To ensure that the mini-batches constructed by the ISF method achieve lower Pad Ratio and Dist ratio, appropriate values for $Q_{v}$ and text $Q_{t}$ need to be determined. The optimal values for $Q_{v}$ and $Q_{t}$ vary across different datasets. In practice, A statistical approach described in Section 5.1 is used to determine these values.
@@ 4.2 Balanced Model Partitioning
4.3 Balanced Adaptive Re-Computation
We find that heterogeneous architectures have different memory requirements. For example, the vision model in InternVL-Chat-1.5 requires more GPU memory than the language model under the same computational load. Therefore, it is necessary to analyze the memory requirements of each layer in the vision and language models individually and adaptively determine the optimal re-computation strategy for each layer.
Balanced Adaptive Re-Computation Strategy. In this context, $Q_{v}$ and $Q_{t}$ represent the inputs for Vision Transformer (VIT) and Large Language Model (LLM), respectively. $M_{r}$ denotes the remaining GPU memory at the current stage, while $M_{t}$ and $M_{v}$ indicate the GPU memory saved by each transformer layer of the LLM and VIT when enabling re-computation.
5 Experiments
In this section, the models and datasets are introduced. Then, we demonstrate the acceleration compared to the current state-of-the-art VLMs. Subsequently, a detailed comparison of each component proposed in our method is presented, highlighting its specific contribution to training acceleration. Finally, extensive experimental analysis is conducted.
@@ 5.1 Experimental setup
5.2 Implementation Details
How to get $Q_{v}$ and $Q_{t}$. We determine $Q_{v}$ and $Q_{t}$ using dataset statistics. The total text token lengths and image count are used to compute the average tokens per image. $Q_{t}$ is set to the longest text token length, and the text-to-image ratio determines $Q_{v}$. It is challenging to maintain an exact number of images and text length, so we relax these conditions to allow for approximation. For images, $Q^{\prime}_{v}=Q_{v}$, and for text, $Q^{\prime}_{t}=Q_{t}-128$ based on results in Table 6. In the InternVL-Chat-1.2 dataset, $Q_{t}=4K$, $Q_{v}=9$, with each image processed into 1K tokens for VIT.
Table 6: Ablation results of $Q_{v},Q_{t}$
@@ 5.2 Implementation Details
Sampling Overhead. The sampling overhead is minimal. For instance, with 1.2 million samples, sampling can be completed in just a few tens of seconds, adding negligible overhead to the overall training process. The time complexity of ISF is $\mathcal{O}(C\cdot(N+M))$, where $N$ is the total number of samples, $M$ is the number of samples per pack, and $C$ is the number of iterations.
Partition Overhead. Establishing the anchor requires profiling layer-wise forward time and activation values, which takes only 5 steps and incurs minimal overhead. While the anchor may not be optimal, we observe empirically that the true optimum typically lies within $r\leq 3$ of it.
Optimal Solution Overhead. To refine the anchor, we measure the speed of the top 10–15 nearby partitions (as in Section 4.2), each requiring just 3 steps. This adds only a minor cost relative to the full training process.
@@ 5.3 Main Results
5.4 Ablation Analysis
In this section, ablation experiments on each component of our method are conducted, using InternVL-Chat-1.5 as the baseline model with a 3D parallel Megatron-DeepSpeed backend. Table 7 illustrates the impact of each component. The baseline model experiences a considerable slowdown in training speed due to computational imbalance, necessitating a total of 61.8 GPU days. By achieving data balance, GPU days are reduced to 51.9. Data balance allows us to achieve a more balanced model partition, reducing the training time. Finally, optimizing memory with an adaptive re-computation strategy reduces GPU days to 21.3. These results demonstrate that a holistic balance encompassing data, model, and memory is crucial for efficient VLM training. Below, we provide a detailed analysis of each component.
The Importance of Data Balance: In Table 3, we investigate the importance of data balance in large-scale distributed training by comparing four methods: (1) Baseline: Randomly combining data into a mini-batch with padding aligned to the longest input within mini-batches, (2) Length-Group: Combining samples with similar text and image sizes into a mini-batch to minimize padding within mini-batches. (3) Device-Group: Grouping samples with similar input sizes across devices to minimize idle times. (4) Balanced Dynamic Mini-batch: Using ISF to construct balanced mini-batches within mini-batches and across devices.
The Importance of Model Balance: Table 4 examines balanced model partitioning, focusing on partition strategies for pipeline parallelism. For LLM training, common methods include (1) parameter-based and (2) layer-based, (3) profile-based methods such as DreamPipe (Narayanan et al., 2019), which estimate the computation time for each layer, and use this information to partition the model effectively. Additionally, (4) our search-based Balanced Model Partition method finds the optimal partition strategy from a set of candidates. As shown in Table 4, (1) Parameter-based and (2) layer-based methods split the model’s parameters or layers across devices, achieving low variation in VAR(param) and VAR(num_layer). However, they still show high variation in forward time VAR(fwd_time), leading to computational inefficiencies in the pipeline. (3) The profile-based method ensures the optimal VAR(fwd_time). However, this partitioning occurs before the vision model’s token sub-sampling operation, increasing communication overhead and affecting training speed. (4) Our proposed BMP method explores a high-quality partition strategy space to identify the optimal strategy, achieving the best results in 29.0 GPU days.
The Importance of Memory Balance: In Table 5, we examine the significance of memory balance. In the baseline model, varying input sizes for vision (4K–20K tokens) and language (1K–16K tokens) lead to varying GPU memory usage. Despite aggressive re-computation, the remaining memory can drop to 7.3 GB. ISF and BMP improve training speed by controlling computational load across devices. However, memory demands still varied, e.g., GPUs 1 and 2 having more remaining memory. ISF further speeds up training by adjusting the re-computation strategy to fully utilize the remaining memory, reducing GPU days to 21.3.
5.5 Generalization Capability
We study the generalization capability of our method from multiple aspects: (1) Different Datasets like LLava-665K, InternVL-1.2M, and LCS558K in Appendix B; (2) Different Models: we verify our methods on different vision models like EVA-CLIP (Sun et al., 2023) and different language models like Llama3 (AI, 2024), Yi-34B (NousResearch, 2023), and the Qwen1.5-110B (Bai et al., 2023a), as detailed in Appendix C; (3) Different High-Resolution Setting: Under various settings, we achieved a speedup of approximately 2.0$\times$, as demonstrated in Appendix D; (4) Different Tasks: Besides SFT tasks, pretraining tasks are also tested, as shown in Appendix E, and we observed consistent improvements across all settings; (5) Different Image Resolutions: As shown in Appendix F, our method consistently delivers a highly satisfactory acceleration effect with different input image resolutions; (6) Convergence of ISF: We show that ISF can converge in a few steps on other datasets, in Appendix G; (7) Different Model-series like LLava-1.6 in Appendix H; (8) Pre-Processing Strategy: Qwen2-VL (team, 2024) is a recent, highly regarded open-source project that provides strong support for dynamic image input. We adopt the pre-processing strategy of Qwen2-VL to validate our method in Appendix I; (9) Different Hardware like H100, A100 in Appendix J; (10) Larger-Scale Results. To validate our method, we conduct experiments with larger models and 512 GPUs. Results are provided in Appendix K. These results underscore the effectiveness and robustness of our method across a wide range of datasets, models, and tasks.
6 Conclusion
Acknowledgements
This work was supported by the National Key Research and Development Program of China under Grant No. 2023YFB4405102. It was also supported in part by the Natural Science Foundation of Hunan Province under Grant No. 2024JJ6525. We would like to thank Kaihuan Liang, Shihao Bai, and Shuo Wu for their support on this work.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Training-Pipeline of our balanced method
A.1 Training-Pipeline
Appendix B Results on different datasets
We consistently achieved a low Dist Ratio across the LLava-665K, InternVL-1.2M, and LCS558K datasets, as demonstrated in Table 8. Additionally, our approach significantly enhanced training speed.
Table 8: Results on different datasets
Dataset Dist Ratio GPU Days
VIT LLM
LLava-665K 0.02 0.145 43.3$\rightarrow$ 12.4 (3.5$\times$)
InternVL-1.2M 0.02 0.14 61.8$\rightarrow$ 21.3 (2.9$\times$)
LCS-558K 0.001 0.029 23.8$\rightarrow$ 7.5 ( 3.2$\times$)
Appendix C Results on different Model Size
We test various combinations of vision and language models. As shown in Table 9, our approach significantly reduces the required GPU days for model training, achieving nearly a 2x speedup across models of various sizes.
@@ Appendix C Results on different Model Size
Appendix D Results on Different Dynamic High-Resolution Settings
To validate the effectiveness of our method, we test it under various high-resolution settings. Our approach consistently demonstrates low Dist Ratio and strong acceleration across all configurations, as shown in Table 10, significantly improving training speed under different settings.
@@ Appendix D Results on Different Dynamic High-Resolution Settings
Appendix E Results on Pretrain Setting
We evaluate our method in other tasks, such as pre-training tasks.. In Pre-training, we train both the Vision Transformer (ViT) and MLP components for models ranging from 6B to 20B. However, for larger models, such as 6B-34B and 6B-70B, we focus solely on training the MLP component. Across all configurations, we observe consistent performance improvements shown in Table 11, particularly with the largest model, where GPU days are significantly reduced from 16.8 to 9.6, demonstrating enhanced training efficiency.
@@ Appendix E Results on Pretrain Setting
Appendix F Results on Different Resolutions
We further test our method with different image resolution inputs. As shown in Table 12, our method consistently delivers low Dist Ratio and highly satisfactory acceleration results across varying image resolutions, demonstrating its effectiveness in improving training efficiency.
@@ Appendix F Results on Different Resolutions
Appendix G Convergence of ISF.
Appendix H Results on Open-source LLava-1.6
We also validate our method using another popular open-source model, LLava-1.6, with the DeepSpeed backend, as shown in Table 13. For the DeepSpeed backend, we employ only our balanced dynamic mini-batch strategy. In the case of the open-source LLava model, while its ViT component is relatively small and the imbalance occurs primarily at the data level, we still achieved a notable overall speedup. Although the speedup ratio is smaller compared to other models, our method delivered a 30% improvement in performance.
@@
Appendix I Results on Qwen2-VL Pre-Processing Strategy
Qwen2-VL is a recent, highly regarded open-source project that provides strong support for dynamic image input. Consequently, we adopt the pre-processing strategy of Qwen2-VL to validate our method. As shown in Table  14, our approach demonstrates a substantial acceleration effect (approximately 1.9x) when applied to the Qwen2-VL strategy, significantly reducing both the padding ratio and dist ratio.
@@ Appendix I Results on Qwen2-VL Pre-Processing Strategy
Appendix J Different Hardware Results
We test our method on various hardware platforms with different GPUs (e.g., A100, H100) and network bandwidths. The experiments in Table  15 confirmed consistent performance improvements across all platforms.
Table 15: Results on Different Hardware. IB indicates network bandwidths.
@@
Appendix K Large-Scale Results
To validate the effectiveness of our method, we conduct a study using larger-scale models and a greater number of GPUs. As shown in the Table 16, our method achieves a speedup ratio exceeding 2.0 across varying GPU configurations. Moreover, the results demonstrate that our approach maintains a more favorable linear speedup (85% $\rightarrow$ 95%) as GPUs increase.